On January 2, 2018, All Good Company purchased 5,000 shares of the stock of Big Bad Company, and DID obtain significant influence. The investment is intended as a long-term investment. The stock was purchased for $13 per share, and represents a 30% ownership stake. Big Bad Company made $350,000 of net income in 2018, and paid dividends to All Good Company of $35,000 on December 15, 2018. Big Bad Company's stock was trading on the open market for $14 per share at the end of the year. Use this information to determine the book value of the investment that should be reported at year end by All Good Company. Round to nearest whole dollar.
Answer #1

B 2 Calculate ending balance of investment 3 4 Initial investment 65000) 5 Add: Net income 105000 6 Less: Dividend 35000 7 En2 Calculate ending balance of investment 3 4 Initial investment =5000*13 5 Add: Net income =350000*30% 6 Less: Dividend 35000